# Composite Indicator-Guided Infilling Sampling for Expensive Multi-Objective Optimization

### v1

@@ Abstract
In expensive multi-objective optimization, where the evaluation budget is strictly limited, selecting promising candidate solutions for expensive fitness evaluations is critical for accelerating convergence and improving algorithmic performance. However, designing an optimization strategy that effectively balances convergence, diversity, and distribution remains a challenge. To tackle this issue, we propose a composite indicator-based evolutionary algorithm (CI-EMO) for expensive multi-objective optimization. In each generation of the optimization process, CI-EMO first employs NSGA-III to explore the solution space based on fitness values predicted by surrogate models, generating a candidate population. Subsequently, we design a novel composite performance indicator to guide the selection of candidates for real fitness evaluation. This indicator simultaneously considers convergence, diversity, and distribution to improve the efficiency of identifying promising candidate solutions, which significantly improves algorithm performance. The composite indicator-based candidate selection strategy is easy to achieve and computes efficiency. Component analysis experiments confirm the effectiveness of each element in the composite performance indicator. Comparative experiments on benchmark problems demonstrate that the proposed algorithm outperforms five state-of-the-art expensive multi-objective optimization algorithms.
Index Terms:
@@
I Introduction
In recent decades, evolutionary algorithms (EAs) have demonstrated remarkable effectiveness in addressing multi- and many-objective optimization problems [1, 2, 3]. These population-based metaheuristic approaches are highly effective at generating diverse and well-converged solutions within a single optimization run [4, 5]. Optimization problems with two or three objectives are typically referred to as multi-objective optimization, while those involving more than three objectives are categorized as many-objective optimization [6]. A wide variety of multi- and many-objective optimization algorithms have been introduced in the literature, which are commonly grouped into three categories: dominance-based [7], decomposition-based [8], and indicator-based approaches [1].
However, in real-world engineering applications, the objective evaluations of many optimization problems often lack explicit mathematical expressions. Instead, these evaluations typically rely on software simulations or physical experiments, both of which can be highly time-consuming [9, 10], such as airfoil design [11], lightweight design of vehicle [12], and network architecture search [13]. These kinds of problems are named expensive multi/many-objective optimization problems (EMOPs/EMaOPs). In these problems, objective function evaluation generally is computationally time-consuming, which hinders the application of traditional multi-objective evolutionary algorithms (MOEAs) [14]. For this reason, surrogate models are widely employed in EMOPs to reduce computational costs by using efficient machine learning methods to approximate the real objective values [15, 16]. Surrogate-assisted evolutionary algorithms (SAEAs) leverage previously accumulated data to build surrogate models. During the optimization process, candidate solutions are initially evaluated using these trained models, and only the most promising ones undergo expensive real fitness evaluations (FEs). This approach effectively reduces the number of expensive evaluations, thereby saving computational time [17, 18].
@@ I Introduction
Fig. 1 shows the flowchart of general surrogate-assisted MOEAs. In solving EMOPs, there are usually only a few hundred expensive FEs available [19, 20]. Therefore, all data evaluated by the real fitness function are generally archived in a database. After initializing the database, optimization consists of three main parts:
@@
    Build/update surrogate models: Surrogate models are built based on data in the database. The common surrogate models include Gaussian Process model (GP, also named Kriging model) [21, 22], radial basis function network (RBFN) [23], and support vector machine (SVM) [24] et al. During the optimization process, the surrogate models are updated in response to changes in the database.
@@
By instantiating specific parts of the generic framework, numerous surrogate-assisted MOEAs have been developed to address expensive multi-objective optimization problems [16, 25, 15]. Since the real fitness evaluation budget is very limited, it is more advantageous to allocate fitness evaluations to the most promising candidate solutions. This raises a key challenge: how to efficiently identify promising solutions during evolutionary search and candidate selection.
For this issue, some representative works have been presented recently. On one hand, some multiobjective efficient global optimization (MOEGO) methods have been proposed based on Bayesian optimization, in which an appropriate expected improvement (EI) function is designed to sample candidates for expensive evaluation [26, 27]. Zhao et al. present an expected direction-based hypervolume improvement to select candidates. After that, an improved method selects a subset of query points based on the lower bound of R2-based EI from the candidates which are generated by the Tchebycheff decomposition paradigm [28]. In these ways, some solutions with uncertainty increments can be retained to improve the accuracy of the alternative model. However, due to the limited number of training samples available for surrogate models, the estimation of uncertainty information can often be inaccurate, which adversely impacts the ability to preserve promising solutions [16]. On the other hand, Chugh et al. proposed selecting one individual from each cluster after the Kriging-assisted reference vector-guided EA search, based on either the angle penalized distance or maximum uncertainty [29]. Song et al. divide optimization into three states after generating candidates using Two_Arch2 [30], and use their respective sampling strategies (convergence, diversity, and uncertainty sampling) to select candidate solutions, respectively [31]. In addition to methods that use sampling strategies for a single purpose, Qin et al. proposed a performance indicator that considers both convergence and diversity properties of candidate solutions by two types of distance calculation [32]. However, current approaches lack consideration of the distribution of solutions. Overall, the above candidate solution selection methods still do not balance convergence, diversity, and distribution in the search process. Distribution means the uniform distribution of the solution set on the Pareto front. To differentiate from the distribution, in this paper, the term "diversity" emphasizes the spread of solution, that is, the expansion of the solution boundary, and whether the extreme areas of the target space are fully explored [1].
@@
    The composite indicator is incorporated into surrogate-assisted MOEAs for expensive multi-objective optimization. At each iteration, CI-EMO first employs GP-assisted NSGA-III [33] as the optimizer to generate the candidate population. The composite indicator is then used to select the most promising candidate for real fitness evaluation, achieving efficient sampling and improving the solution set’s distribution. Not only is the algorithm easy to reproduce, but experiments have shown that it performs significantly better than the state-of-the-art expensive multi-objective optimization algorithms.
The rest parts of this paper are organized as follows. Section II introduces the background of this work. In Section III, we introduce our proposed composite indicator and CI-EMO algorithm. In Section IV, a series of experiments are designed to validate the performance of our proposed method. Finally, Section V offers a conclusion.
II BACKGROUND
II-A Problem Formulation
@@
where $\boldsymbol{x}$ represents the decision variables, and $\Omega\subseteq\mathbb{R}^{d}$ defines the feasible decision space. The function $F:\Omega\to\mathbb{R}^{m}$ is a vector function containing $m$ continuous objectives. In the context of this study, we focus on cases where evaluating the functions is computationally expensive but can be done in parallel, for example, using multiple processors or machines [4].
Let $\boldsymbol{x}_{a}$ and $\boldsymbol{x}_{b}$ represent two vectors in the decision space. Vector $\boldsymbol{x}_{a}$ is said to dominate $\boldsymbol{x}_{b}$ (denoted as $\boldsymbol{x}_{a}\prec\boldsymbol{x}_{b}$) if and only if $f_{i}(\boldsymbol{x}_{a})\leq f_{i}(\boldsymbol{x}_{b})$ holds for all $i=1,\dots,m$, with at least one objective index $j$ satisfying $f_{j}(\boldsymbol{x}_{a})<f_{j}(\boldsymbol{x}_{b})$. A solution $\boldsymbol{x}^{*}\in\Omega$ is deemed Pareto optimal if no other solution dominates it. The corresponding objective vector, $F(\boldsymbol{x}^{*})$, is referred to as the Pareto optimal objective vector. The collection of all Pareto optimal solutions constitutes the Pareto set (PS), while the set of all associated Pareto optimal objective vectors defines the Pareto front (PF) [34].
II-B Gaussian Process Model
Gaussian Process model is a machine learning technique used for spatial interpolation, offering predictions of objective values for solutions along with an estimate of the uncertainty in these predictions [21].
@@ II-B Gaussian Process Model
III PROPOSED ALGORITHM
III-A CI-EMO Framework
The framework of CI-EMO is illustrated in Fig. 2. The CI-EMO algorithm begins by initializing a database through Latin Hypercube Sampling (LHS) [35]. Each generation of the algorithm consists of two key stages: (1) Candidates generation and (2) Composite indicator-based selection. During the optimization process, CI-EMO first constructs Gaussian Process surrogate models for each objective based on the database. It then leverages the surrogate-assisted NSGA-III [33] (SA-NSGA-III) algorithm to explore the search space over a predefined number of generations, resulting in a candidate population. Subsequently, CI-EMO employs a composite indicator that synthesizes three distinct metrics, i.e., distribution, diversity, and convergence, to select the most promising sample point from the candidate population for expensive fitness evaluation. The evaluated data point is then incorporated into the database, facilitating the refinement of surrogate models for subsequent generations.
@@ 8       $\mathcal{D}\leftarrow\mathcal{D}\cup\left\{\boldsymbol{x},F\left(\boldsymbol{x}\right)\right\}$;
10 end while
11Return: All nondominated solutions in $\mathcal{D}$;
@@
III-B Generate candidates
To enhance the convergence rate of the algorithm and effectively identify promising sample points, CI-EMO leverages SA-NSGA-III during the optimization process to explore the solution space and generate a candidate population for subsequent selection. The pseudocode of SA-NSGA-III is presented in Algorithm 2. Initially, the input consists of the database $\mathcal{D}$ and surrogate models $\mathcal{M}$. The algorithm generates a set of reference vectors, denoted as $V$, and computes the ideal point $Z$ based on the database. Using the environmental selection mechanism of NSGA-III [33], a non-dominated population $P$ is obtained. The subsequent search procedure follows the NSGA-III framework, where the fitness values of offspring are predicted by surrogate models. Specifically, offspring $O$ are generated through crossover and mutation operations, and then merged with population $P$. Once the maximum number of iterations, $T_{max}$, is reached, the algorithm outputs the candidate population $P^{*}$.
@@ 8 $\boldsymbol{x}\leftarrow$ Select the solution with maximum $\mathbf{CI}$ value in $P^{*}$;
III-C Composite indicator-based selection
The combination of indicators is an important method to generate new selection mechanisms with different requirements [1, 36]. To improve the sampling effect for the expensive multi-objective problem, CI-EMO combines three sampling metrics into a composite indicator. The three metrics consider distribution, diversity, and convergence, respectively. Algorithm 3 presents the pseudocode for the composite indicator selection used in CI-EMO. The composite performance of each candidate solution in the $P^{*}$ will be evaluated in terms of distribution, diversity, and convergence metrics. Note that, the three component indicators are calculated only based on $P^{*}$ and $\mathcal{D}$. As shown in algorithm 3, before calculating the distribution indicator, the algorithm performs non-dominated sorting on $\mathcal{D}$ then obtains the non-dominated solution set $\mathcal{D}_{nd}$. The distribution indicator is calculated based on the original objective value. However, before calculating diversity and convergence indicators, the normalized objective function of solutions $P^{*}$ and $\mathcal{D}$ is calculated, respectively, as follows:
@@ III-C Composite indicator-based selection
    Diversity indicator ($\mathbf{I_{2}}$): The diversity of a candidate solution $\boldsymbol{x}$ is defined based on the crowdedness surrounding the solution. It also is termed as MaxMin diversity indicator [37]. Compared to distribution, the indicator emphasizes the spread of solution, that is, the expansion of the solution boundary, and whether the extreme areas of the target space are fully explored [1]. It is quantified by computing the distance in the objective space to its nearest neighbor $\tilde{{\boldsymbol{F}}}\left(\boldsymbol{y}\right)$, as expressed by the Eq. 7. $\mathbf{I}_{2}$ is calculated by normalizing $d_{c}$.
@@
IV EXPERIMENTAL STUDIES
To verify the performance of the proposed CI-EMO, comparative studies are conducted by comparing CI-EMO with five state-of-the-art surrogate-assisted MOEAs on multi- and many-objective benchmarks. Compared algorithms include K-RVEA [29], KTA2 [31], EMMOEA [32], DirHV-EGO [38], and R2/D-EGO [28]. Then several experiments were conducted to analyze the effectiveness of CI-EMO, including component analysis of composite indicator, ablation studies, and sampling number analysis. All algorithms are implemented in MATLAB using the PlatEMO platform [39]. To guarantee consistency and reliability, all experiments are conducted on identical hardware specifications (CPU: Intel(R) Core(TM) i9-14900KF, Memory: 128GB).
IV-A Experimental Settings

### v2

@@ Abstract
In expensive multi-objective optimization, where the evaluation budget is strictly limited, selecting promising candidate solutions for expensive fitness evaluations is critical for accelerating convergence and improving algorithmic performance. However, designing an optimization strategy that effectively balances convergence, diversity, and distribution remains a challenge. To tackle this issue, we propose a composite indicator-based evolutionary algorithm (CI-EMO) for expensive multi-objective optimization. In each generation of the optimization process, CI-EMO first employs NSGA-III to explore the solution space based on fitness values predicted by surrogate models, generating a candidate population. Subsequently, we design a novel composite performance indicator to guide the selection of candidates for real fitness evaluation. This indicator simultaneously considers convergence, diversity, and distribution to improve the efficiency of identifying promising candidate solutions, which significantly improves algorithm performance. The composite indicator-based candidate selection strategy is easy to achieve and computes efficiency. Component analysis experiments confirm the effectiveness of each element in the composite performance indicator. Comparative experiments on three benchmark test sets and real-world problems demonstrate that the proposed algorithm outperforms five state-of-the-art expensive multi-objective optimization algorithms.
keywords:
@@
journal: Elsevier
1 Introduction
In recent decades, evolutionary algorithms (EAs) have demonstrated remarkable effectiveness in addressing multi- and many-objective optimization problems (Falcón-Cardona and Coello, 2020; Zhang et al., 2024a; Rostamian et al., 2025). These population-based metaheuristic approaches are highly effective at generating diverse and well-converged solutions within a single optimization run (Liu et al., 2023b; Zhang et al., 2025). Optimization problems with two or three objectives are typically referred to as multi-objective optimization, while those involving more than three objectives are categorized as many-objective optimization (Li et al., 2015). A wide variety of multi- and many-objective optimization algorithms have been introduced in the literature, which are commonly grouped into three categories: dominance-based (Zhou et al., 2024), decomposition-based (Li, 2024), and indicator-based approaches (Falcón-Cardona and Coello, 2020).
However, in real-world engineering applications, the objective evaluations of many optimization problems often lack explicit mathematical expressions. Instead, these evaluations typically rely on software simulations or physical experiments, both of which can be highly time-consuming (Zhen et al., 2021; Zhai et al., 2024), such as airfoil design (Dai et al., 2024), lightweight design of vehicle (Li et al., 2024), and network architecture search (Li et al., 2025). These kinds of problems are named expensive multi/many-objective optimization problems (EMOPs/EMaOPs). In these problems, objective function evaluation generally is computationally time-consuming, which hinders the application of traditional multi-objective evolutionary algorithms (MOEAs) (Li et al., 2022). For this reason, surrogate models are widely employed in EMOPs to reduce computational costs by using efficient machine learning methods to approximate the real objective values (Wei et al., 2025; Zhang et al., 2024b). Surrogate-assisted evolutionary algorithms (SAEAs) leverage previously accumulated data to build surrogate models. During the optimization process, candidate solutions are initially evaluated using these trained models, and only the most promising ones undergo expensive real fitness evaluations (FEs). This approach effectively reduces the number of expensive evaluations, thereby saving computational time (Jin et al., 2018; He et al., 2023).
@@ 1 Introduction
Fig. 1 shows the flowchart of general surrogate-assisted MOEAs. In solving EMOPs, there are usually only a few hundred expensive FEs available (Zhang et al., 2024b; Yang et al., 2024). Therefore, all data evaluated by the real fitness function are generally archived in a database. After initializing the database, optimization consists of three main parts:
@@
    Build/update surrogate models: Surrogate models are built based on data in the database. The common surrogate models include Gaussian Process model (GP, also named Kriging model) (Schulz et al., 2018; Wang et al., 2025), radial basis function network (RBFN) (Chen et al., 2023), and support vector machine (SVM) (Sonoda and Nakata, 2022) et al. During the optimization process, the surrogate models are updated in response to changes in the database.
@@
By instantiating specific parts of the generic framework, numerous surrogate-assisted MOEAs have been developed to address expensive multi-objective optimization problems (Zhang et al., 2024b; Liu et al., 2023a; Wei et al., 2025). Since the real fitness evaluation budget is very limited, it is more advantageous to allocate fitness evaluations to the most promising candidate solutions. This raises a key challenge: how to efficiently identify promising solutions during evolutionary search and candidate selection.
For this issue, some representative works have been presented recently. On one hand, some multiobjective efficient global optimization (MOEGO) methods have been proposed based on Bayesian optimization, in which an appropriate expected improvement (EI) function is designed to sample candidates for expensive evaluation (Xu et al., 2025; Zhan and Xing, 2020). Zhao et al. present an expected direction-based hypervolume improvement to select candidates. After that, an improved method selects a subset of query points based on the lower bound of R2-based EI from the candidates which are generated by the Tchebycheff decomposition paradigm (Zhao et al., 2024). In these ways, some solutions with uncertainty increments can be retained to improve the accuracy of the alternative model. However, due to the limited number of training samples available for surrogate models, the estimation of uncertainty information can often be inaccurate, which adversely impacts the ability to preserve promising solutions (Zhang et al., 2024b). On the other hand, Chugh et al. proposed selecting one individual from each cluster after the Kriging-assisted reference vector-guided EA search, based on either the angle penalized distance or maximum uncertainty (Chugh et al., 2016). Song et al. divide optimization into three states after generating candidates using Two_Arch2 (Wang et al., 2014), and use their respective sampling strategies (convergence, diversity, and uncertainty sampling) to select candidate solutions, respectively (Song et al., 2021). In addition to methods that use sampling strategies for a single purpose, Qin et al. proposed a performance indicator that considers both convergence and diversity properties of candidate solutions by two types of distance calculation (Qin et al., 2023). However, current approaches lack consideration of the distribution of solutions. Overall, the above candidate solution selection methods still do not balance convergence, diversity, and distribution in the search process. Distribution means the uniform distribution of the solution set on the Pareto front. To differentiate from the distribution, in this paper, the term "diversity" emphasizes the spread of solution, that is, the expansion of the solution boundary, and whether the extreme areas of the target space are fully explored (Falcón-Cardona and Coello, 2020).
@@
    Experiments validate that the proposed composite indicator is better than single one. The composite indicator-guided evolutionary algorithm is a promising direction for candidate selection in SAEAs.
  3. 3.
    The composite indicator is incorporated into surrogate-assisted MOEAs for expensive multi-objective optimization. At each iteration, CI-EMO first employs GP-assisted NSGA-III (Deb and Jain, 2013) as the optimizer to generate the candidate population. The composite indicator is then used to select the most promising candidate for real fitness evaluation, achieving efficient sampling and improving the solution set’s distribution.
The rest parts of this paper are organized as follows. Section 2 introduces the background of this work. In Section 3, we introduce our proposed composite indicator and CI-EMO algorithm. In Section 4, a series of experiments are designed to validate the performance of our proposed method. Finally, Section 5 offers a conclusion.
2 BACKGROUND
2.1 Problem Formulation
@@
where $\boldsymbol{x}$ represents the decision variables, and $\Omega\subseteq\mathbb{R}^{d}$ defines the feasible decision space. The function $F:\Omega\to\mathbb{R}^{m}$ is a vector function containing $m$ continuous objectives. In the context of this study, we focus on cases where evaluating the functions is computationally expensive but can be done in parallel, for example, using multiple processors or machines (Liu et al., 2023b).
Let $\boldsymbol{x}_{a}$ and $\boldsymbol{x}_{b}$ represent two vectors in the decision space. Vector $\boldsymbol{x}_{a}$ is said to dominate $\boldsymbol{x}_{b}$ (denoted as $\boldsymbol{x}_{a}\prec\boldsymbol{x}_{b}$) if and only if $f_{i}(\boldsymbol{x}_{a})\leq f_{i}(\boldsymbol{x}_{b})$ holds for all $i=1,\dots,m$, with at least one objective index $j$ satisfying $f_{j}(\boldsymbol{x}_{a})<f_{j}(\boldsymbol{x}_{b})$. A solution $\boldsymbol{x}^{*}\in\Omega$ is deemed Pareto optimal if no other solution dominates it. The corresponding objective vector, $F(\boldsymbol{x}^{*})$, is referred to as the Pareto optimal objective vector. The collection of all Pareto optimal solutions constitutes the Pareto set (PS), while the set of all associated Pareto optimal objective vectors defines the Pareto front (PF) (Yu, 1974).
2.2 Gaussian Process Model
Gaussian Process model is a machine learning technique used for spatial interpolation, offering predictions of objective values for solutions along with an estimate of the uncertainty in these predictions (Schulz et al., 2018).
@@ 2.2 Gaussian Process Model
3 PROPOSED ALGORITHM
3.1 CI-EMO Framework
The framework of CI-EMO is illustrated in Fig. 2. The CI-EMO algorithm begins by initializing a database through Latin Hypercube Sampling (LHS) (Iman, 2008). Each generation of the algorithm consists of two key stages: (1) Candidates generation and (2) Composite indicator-based selection. During the optimization process, CI-EMO first constructs Gaussian Process surrogate models for each objective based on the database. It then leverages the surrogate-assisted NSGA-III (Deb and Jain, 2013) (SA-NSGA-III) algorithm to explore the search space over a predefined number of generations, resulting in a candidate population. Subsequently, CI-EMO employs a composite indicator that synthesizes three distinct metrics, i.e., distribution, diversity, and convergence, to select the most promising sample point from the candidate population for expensive fitness evaluation. The evaluated data point is then incorporated into the database, facilitating the refinement of surrogate models for subsequent generations.
@@ 8       $\mathcal{D}\leftarrow\mathcal{D}\cup\left\{\boldsymbol{x},F\left(\boldsymbol{x}\right)\right\}$;
10Return: All nondominated solutions in $\mathcal{D}$;
@@
3.2 Generate candidates
To enhance the convergence rate of the algorithm and effectively identify promising sample points, CI-EMO leverages SA-NSGA-III during the optimization process to explore the solution space and generate a candidate population for subsequent selection. The pseudocode of SA-NSGA-III is presented in Algorithm 2. Initially, the input consists of the database $\mathcal{D}$ and surrogate models $\mathcal{M}$. The algorithm generates a set of reference vectors, denoted as $V$, and computes the ideal point $Z$ based on the database. Using the environmental selection mechanism of NSGA-III (Deb and Jain, 2013), a non-dominated population $P$ is obtained. The subsequent search procedure follows the NSGA-III framework, where the fitness values of offspring are predicted by surrogate models. Specifically, offspring $O$ are generated through crossover and mutation operations, and then merged with population $P$. Once the maximum number of iterations, $T_{max}$, is reached, the algorithm outputs the candidate population $P^{*}$.
@@ 8 $\boldsymbol{x}\leftarrow$ Select the solution with maximum $\mathbf{CI}$ value in $P^{*}$;
3.3 Composite indicator-based selection
The combination of indicators is an important method to generate new selection mechanisms with different requirements (Falcón-Cardona and Coello, 2020; Falcón-Cardona et al., 2022). To improve the sampling effect for the expensive multi-objective problem, CI-EMO combines three sampling metrics into a composite indicator. The three metrics consider distribution, diversity, and convergence, respectively. Algorithm 3 presents the pseudocode for the composite indicator selection used in CI-EMO. The composite performance of each candidate solution in the $P^{*}$ will be evaluated in terms of distribution, diversity, and convergence metrics. Note that, the three component indicators are calculated only based on $P^{*}$ and $\mathcal{D}$. As shown in algorithm 3, before calculating the distribution indicator, the algorithm performs non-dominated sorting on $\mathcal{D}$ then obtains the non-dominated solution set $\mathcal{D}_{nd}$. The distribution indicator is calculated based on the original objective value. However, before calculating diversity and convergence indicators, the normalized objective function of solutions $P^{*}$ and $\mathcal{D}$ is calculated, respectively, as follows:
@@ 3.3 Composite indicator-based selection
    Diversity indicator ($\mathbf{I_{2}}$): The diversity of a candidate solution $\boldsymbol{x}$ is defined based on the crowdedness surrounding the solution. It is also termed as MaxMin diversity indicator (Pereverdieva et al., 2025). Compared to distribution, the indicator emphasizes the spread of solution, that is, the expansion of the solution boundary, and whether the extreme areas of the target space are fully explored (Falcón-Cardona and Coello, 2020). It is quantified by computing the distance in the objective space to its nearest neighbor $\tilde{{\boldsymbol{F}}}\left(\boldsymbol{y}\right)$, as expressed by the Eq. 7. $\mathbf{I}_{2}$ is calculated by normalizing $d_{c}$.
@@
4 EXPERIMENTAL STUDIES
To verify the performance of the proposed CI-EMO, comparative studies are conducted by comparing CI-EMO with five state-of-the-art surrogate-assisted MOEAs on multi- and many-objective benchmarks. Compared algorithms include K-RVEA (Chugh et al., 2016), KTA2 (Song et al., 2021), EMMOEA (Qin et al., 2023), DirHV-EGO (Zhao and Zhang, 2023), and R2/D-EGO (Zhao et al., 2024). Then several experiments were conducted to analyze the effectiveness of CI-EMO, including component analysis of composite indicator, ablation studies, and sampling number analysis. All algorithms are implemented in MATLAB using the PlatEMO platform (Tian et al., 2017). To guarantee consistency and reliability, all experiments are conducted on identical hardware specifications (CPU: Intel(R) Core(TM) i9-14900KF, Memory: 128GB).
4.1 Experimental Settings